### v2

# Anchored Learning for On-the-Fly Adaptation — Extended Technical Report

Abstract
This study presents “anchor critics”, a novel strategy for enhancing the robustness of reinforcement learning (RL) agents in crossing the sim-to-real gap. While RL agents can be successfully trained in simulation, they often encounter difficulties such as unpredictability, inefficient power consumption, and operational failures when deployed in real-world scenarios. We identify that naïve fine-tuning approaches lead to catastrophic forgetting, where policies maintain high rewards on frequently encountered states but lose performance on rarer, yet critical scenarios. Our method maximizes multiple Q-values across domains, ensuring high performance in both simulation and reality. Evaluations demonstrate that our approach enables behavior retention in sim-to-sim gymnasium tasks and in sim-to-real scenarios with racing quadrotors, achieving a near-50% reduction in power consumption while maintaining controllable, stable flight. We also contribute SwaNNFlight, an open-source firmware for testing adaptation techniques on real robots. Video: https://youtu.be/WgDFaS5nXrw
I Introduction
Reinforcement Learning (RL) has emerged as a powerful approach for solving complex control problems where traditional techniques may be inadequate. Due to the expensive and potentially risky nature of real-world robot interaction, RL agents are typically trained and evaluated in simulated environments. However, recent work has shown that controllers applied to real systems can exhibit undesirable properties, from minor control issues to severe hardware failures [1, 2, 3]. These problems stem from the challenge of crossing the domain gap between simulated training environments and the real world—the “reality gap”.
While direct on-robot training raises safety and cost concerns, making simulated training attractive, the transition from simulation to reality presents significant challenges[4, 5]. Real-world control scenarios can generate distributions of trajectories where critical scenarios occur only in the tails. This makes the expected discounted sum of rewards, and thus, the associated Q-value, an inadequate measure of agent performance across important tasks. Recent works have identified that traditional fine-tuning approaches often result in catastrophic forgetting, where policies lose their previously well-behaved performance when adapting to new environments [6, 7]. Indeed, we observed this phenomenon across every benchmark task we tested.
To address these challenges, we propose a method that maximizes multiple Q-values across domains in a multi-objective manner. By leveraging the flexibility of simulated training, we can control the data distribution by evaluating agents on critical scenarios, allowing the Q-value from simulation to serve as effective “anchors” for key performance metrics while adapting to Q-values learned from real-world data. This approach enables robust sim-to-real transfer while preserving essential behaviors learned in simulation.
Our work also addresses the practical challenges of training RL agents online using resource-constrained robotic platforms. Many existing solutions offload computations to ground-station servers, limiting the robots’ autonomy and practicality due to network dependencies and communication constraints [8, 9]. We present SwaNNFlight, an open-source firmware enabling on-board inference with efficient ground station communication for policy updates, making our approach viable for real-world applications requiring live adaptation and real-time control responses.
II Related Work
SwaNNFlight extends Neuroflight, an open-source firmware for NN-based flight control [10, 11]. While Neuroflight demonstrated the potential of embedding NN models into flight control systems, it faced challenges such as high training variance, poor sim-to-real transfer, and limited flexibility in controller updates. We address these issues by tackling the problem of adaptating RL agents between domains and by providing a flexible firmware for live controller updates.
A body of work exists on facing adaptation challenges:
  • $\bullet$
    Model-based methods aim to identify optimal parameters from observations [12, 13, 14], but struggle with accurately modeling complex real-world dynamics.
  • $\bullet$
    Model-free approaches like Meta RL [15, 16, 17] train policies to adapt to different environments, but are limited by the fidelity of simulated environments in representing real-world dynamics shifts.
  • $\bullet$
    Trust Region Policy Optimization (TRPO) [18] and Proximal Policy Optimization (PPO) [19] improve stability by constraining policy changes during optimization. Our anchor critics method shares the goal of preserving existing behavior but focuses on optimizing performance across both simulated and real environments.
  • $\bullet$
    Liquid Neural Networks [20, 21, 22, 23, 24] build adaptability into the network structure itself, showing promise in control tasks with smaller models. However, their effectiveness with out-of-distribution data needs further exploration.
  • $\bullet$
    Some approaches employ RL directly on robotic platforms [25, 26], addressing safety concerns through constrained exploration, though this relies on safety criteria modeling.
  • $\bullet$
    Hebbian plasticity approaches offer adaptive control in soft robotics, addressing morphological damage and improving human-robot interactions. Recent work has applied these techniques to compliant robotics, meta-learning for damage adaptation, and improving handshaking synchronization [27, 28, 29].
It’s important to note that systematic comparison of these approaches is challenging due to the lack of standardized benchmarks for evaluating domain adaptability in reinforcement learning. This underscores the need for more unified evaluation methods in this rapidly evolving field.
III Live Adaptation with Anchor Critics
The main problem our work seeks to address is that of transferring learned control policies between similar domains in real-time. In our main use-case, this is transferring controllers trained in simulation to the real-world, where simulated dynamics may not be sufficiently accurately captured to enable sim-to-real transfer without performance degradation. A common issue of sim-to-real policy transfer is erratic control, which leads to high power consumption and mechanical stress. Our proposed solution attempts to bridge the domain gap during transfer by adapting controllers to the new domain during live operation, aiming to refine and enhance the control response.
Domain Gap
The aforementioned ‘domain gap’ is characterized by a change in the underlying dynamics between two domains. Attempting to simply adapt policies learned in simulation to new data derived from the new dynamics was observed to quickly lead to catastrophic forgetting where ‘good’ behavior learned in simulation would be forgotten, causing the policies to become unstable and unpredictable. [6, 7] When new observations are not consistent with simulated observations, unpredictably aberrant behavior is a safety risk. Unlike in simulation, it is not possible in real-life to safely initialize the control system in arbitrary and diverse conditions. Experience in the new domain needs to start being gathered from a ‘safe’ region and boundaries of safety can be expanded as the policy adapts to the new domain. However, this induces a bias in the adapted policy towards those initial ‘safe’ experiences – and while some degree of bias is beneficial for the policy to adapt itself, balancing the effects of the new experience against the beneficial experiences gained in simulation can be tricky.
A Case Against Mixing Experience Buffers During Adaptation
It may seem reasonable, as a first-pass solution, to adapt policies by mixing simulated and real experience during fine-tuning. However, this can be problematic in two critical ways. (i) It breaks the Markov assumption by introducing unobserved hidden variables relating to which dynamical environment any given sample of experience may have come from, thus introducing uncertainty and inconsistency to the relationship between states, actions, and rewards – where consistency is critical for effective RL training. (ii) It would be difficult to ensure a balance in the diversity of the two disparate datasets. This could introduce a reward skew, as discussed further in Section III-A. Instead, we suggest that separating the value functions that drive policy optimization between domains would enable weighing how much policies ’care’ about one domain over the other. Thus allowing a tunable trade-off of stability over adaptability and vice versa.
Anchor-Critics
To mitigate the effects of domain transfer and adapting to a new domain live during active deployment, we introduce Anchor-Critics (more details in Section III-C). Anchors are an augmentation to typical Actor-Critic optimization, where we keep the value function trained on the initial (e.g. simulated) domain during policy fine-tuning to ‘anchor’ the policy to its previously learned behavior, and train a secondary, independent critic to reflect value estimation on the target domain (e.g. real-world). The balance between the anchor and the new critic is tunable, and consequently, how much a policy is allowed to deviate from previously learned behavior during fine-tuning for adaptation is also tunable. We are able to take some of the guess-work out of experience balancing and ensure better stability by limiting policy drift. In the following sections, we will demonstrate the utility of anchors across a range of benchmark RL environments as well as for the motivating problem of live policy adaptation on a quadrotor drone.
III-A Data Distribution Impact on Policy Optimization
Multi-objective optimization is significantly impacted by the coverage of the data upon which solutions are optimized. Consider, for example, a two-objective policy optimization which aims to simultaneously minimize actuation (analogous to power consumption), $\phi_{p}$, and goal-tracking error, $\phi_{e}$, such that the optimal policy, $\pi^{*}$ is given by:
$\pi^{*}=\operatorname*{arg\,min}_{\pi}\left(f(\phi_{p}(s),\phi_{e}(s))\right)\ \ \forall s\in S$ (1)
for all states $s$ in the set of valid states and with objectives composed by some objective composition function $f$. With a sufficiently diverse set of training examples, $\hat{S}\approx S$, good coverage allows a policy optimized over $\hat{S}$ would be a reasonable approximation of the optimal policy. However, if coverage is limited to a limited region of the state-space such that one of the criteria becomes degenerate this could drive policies into trivial optima. For example, if a majority of the data involves a drone hovering in place, where $\phi_{e}\approx 0$), the optimal actuation would be no actuation, but this would also correspond to an abdication of control. This trivial solution is at odds with the requirements of controlability. A formalized example of this issue is given in Section III-B.
In the specific example of high-performance racing drone control, where swift aerobatic response is critical, the safest control scenario in which a policy trained in simulation can be deployed is to first tune a policy for stable flight in slower maneuvers. This biases the fine-tuning data capture towards low actuation as new data is collected, even with past simulated experiences retained in a replay buffer as their proportion diminishes, and causes the policy to forget more agile control. As actor-critic methods make use of replay buffers, the issue of forgetting appears in various forms accross multiple tasks we explore on a real quadrotor in Section VI-B and on benchmark control tasks in Section IV-C.
III-B Illustrating Reward Skew
Consider a reward function $R(s,a)$ linearly composed of two components, both bounded between 0 and 1 such as
$R(s,a)=w_{p}R_{p}(s,a)+w_{d}R_{d}(s,a),$ (2)
where $R_{p}$ is the reward component for power efficiency, $R_{d}$ is the reward component for distance to target, and $w_{p},w_{d}$ are their respective weights.
Let us separately define $R_{p}$ and $R_{d}$ as follows:
$R_{p}(s,a)=1-\frac{\|a\|^{2}}{\|a_{\text{max}}\|^{2}},$ (3)
where $a$ is the action vector and $a_{\text{max}}$ is the maximum possible action vector. Smaller actions are considered more power efficient in this scenario. Next, we define
$R_{d}(s,a)=1-\frac{\|s-s_{\text{setpoint}}\|}{\|s_{\text{max}}-s_{\text{setpoint}}\|},$ (4)
where $\|s-s_{\text{setpoint}}\|$ is the distance between the current state and the setpoint. $R_{d}(s,a)$ therefore represents how near the quadrotor is to the target position and orientation.
By substituting our reward function and separating the expectations, we can define
$Q^{\pi}(s,a)=w_{p}Q_{p}^{\pi}(s,a)+w_{d}Q_{d}^{\pi}(s,a),$ (5)
where $Q_{p}^{\pi}(s,a)$ and $Q_{d}^{\pi}(s,a)$ are the Q-values for power efficiency and distance to target, respectively.
Now, consider two scenarios:
1) The environment distribution is heavily skewed towards states near the target—as is the case when the quadrotor is hovering. In this scenario, $\|s-s_{\text{setpoint}}\|\approx 0$, which implies $R_{d}(s,a)\approx 1$. Consequently we have
$Q_{d}^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot 1|s_{0}=s,a_{0}=a\right]\approx\frac{1}{1-\gamma}.$ (6)
This is the maximum possible value, as $R_{d}$ is always near 1 in this scenario. The final Q-value function becomes
$Q^{\pi}(s,a)\approx w_{p}Q_{p}^{\pi}(s,a)+\frac{w_{d}}{1-\gamma}.$ (7)
In this case, the optimal policy $\pi^{*}$ will be:
$\pi^{*}=\operatorname*{arg\,max}_{\pi}w_{p}V_{p}^{\pi}(s).$ (8)
This policy will focus solely on minimizing power consumption, as the distance to target term is constant and already maximized. Thus, the optimal policy will produce $||a||^{2}=0$, which is to perform no control at all.
2) The environment has a more balanced distribution of states. As such, the optimal policy $\pi^{*}$ will remain
$\pi^{*}=\operatorname*{arg\,max}_{\pi}\left(w_{p}V_{p}^{\pi}(s)+w_{d}V_{d}^{\pi}(s)\right).$ (9)
This optimal policy will balance both power efficiency and trajectory tracking, depending on the relative weights $w_{p}$ and $w_{d}$.
III-C Anchor Critics
Our results suggest that the issue of catastrophic forgetting naturally arises in a wide range of robot controllers when live adaptation with RL is naïvely performed. We observed catastrophic forgetting within only a few updates across various benchmark control problems, underscoring the severity of the issue. In simulation, it is possible to constantly train for a wide-range of maneuvers without risking physical safety. Conversely, many aspects of real robot data may be uncontrollable. Valid live adaptation must incorporate real data while respecting the behavior learned in simulation. We term the domain in which the agent is trained the source domain and the domain in which the agent is adapted the target domain. In order to tackle the issues presented, we propose adapting agents on two main value criteria: (i) the Q-value $Q_{\pi}$ on observed on data from the target and (ii) the Q-value $Q_{\Psi}$ learned on data from the source. We term the latter the ‘anchor’ critic as it serves to anchor the policy behavior to the behavior learned in simulation.
We then treat these Q-values as smooth constraints that must be jointly satisfied. By requiring that step-rewards are normalized—i.e., $R:s\times a\to[0,1]$—we can bound the maximum possible Q-value as follows:
$Q_{\max}=\mathbf{E}_{\tau}[R(\tau)=1]=\sum^{\infty}_{k=0}\gamma^{k}=\frac{1}{1-\gamma}.$ (10)
By dividing the Q-values by $Q_{\max}$ we obtain a normalized Q-value. Treated as a constraint, these normalized Q-values capture how much the current policy satisfies our goals. Since we would like to satisfy multiple Q-values, we then aim to maximize what is known as the product T-norm in Fuzzy Logic. We motivate this choice in Section VII.
The anchored policy optimization goal, given anchor critic $Q_{\Psi}$, extends the classical actor-critic policy loss $J_{\pi_{\theta}}^{\Psi}$ as
$-\mathbb{E}_{s_{r}\sim\mathcal{D}_{\text{T}},s_{s}\sim\mathcal{D}_{\text{S}}}[Q_{\pi_{\theta}}(s_{r},\pi_{\theta}(s_{r}))\cdot Q_{\Psi}(s_{s},\pi_{\theta}(s_{s}))^{w_{\Psi}}]$ (11)
where $\mathcal{D}_{\text{T}}$ and $\mathcal{D}_{\text{S}}$ are the replay buffers for the target and source domains respectively, and $w_{\Psi}$ is the anchor weighting. The amount of weight offered to the anchor dictates the trade-off between satisfying performance objectives in source and target domains.
To keep anchors representative on the source domain, we keep updating $Q_{\Psi}$ during training, following the typical actor-critic update. For this, we use the current adapted policy but only on data sampled from the replay buffer $\mathcal{D}_{\text{S}}$ generated during source domain training. The Q-value $Q_{\pi_{\theta}}$ is conversely trained only on data sampled from the target domain replay buffer $\mathcal{D}_{\text{T}}$.
This simultaneous sampling and optimization approach ensures that the policy maintains a balance between adapting to the target domain data while preserving the knowledge gained from source domain training, effectively mitigating catastrophic forgetting.
(a) DDPG
(b) SAC
(c) TD3
Figure 1: Evolution of the angle ($sin(\theta)$) of an inverted pendulum over time for policies trained with anchor-critics tested on DDPG, SAC and TD3. There are 5 different test runs per goal, and the angle of the pendulum is initialized at random. All 3 anchored algorithms consistently learn to point to 0 rad, thus maintaining a good compromise between our initial goal and our new goal.
IV Anchor Critics in Simulation
In this section, we present a comprehensive evaluation of anchor critics through a series of experiments. First, we demonstrate the general applicability of anchor critics on a toy problem, detailed in Section IV-A. Our initial explorations provide insights into the basic functionality and potential of anchor critics.
Following this, we conduct sim-to-sim transfer evaluations using modified Gymnasium environments in Section IV-C. These modifications are reflect specific adaptation goals, allowing us to showcase the prevalence of catastrophic forgetting in traditional approaches and demonstrate how anchor critics effectively mitigate this issue. We provide anchors as a library and allow the experiments to be reproduced [30].
IV-A Testing Anchors on Inverted Pendulum
In this and the following sections, we evaluate the behavior of agents refined using anchor critics when fine-tuning controllers to establish the effectiveness of the approach in terms of (i) impact on policy stability, (ii) catastrophic forgetting.
(a) $\pi_{\text{\tiny S}}$ on source:
     $\alpha=10^{\circ}$
(b) $\pi_{\text{S}\triangleright\text{T}}$ on target:
     $\alpha=-10^{\circ}$
[…]
     with source anchor
[…]
To test the general viability of anchor critics, we implemented them in three popular contemporary RL algorithms: DDPG [31], SAC [32], and TD3 [33] in Fig. 1 and tested them on a modified inverted pendulum problem. The typical goal of the task is to stabilize a torque-controlled pendulum vertically upright, but we instead aim to hold the pendulum at a specified angle. As shown in Fig. 2, naively fine-tuned, the agents learn exclusively solve the task in the source or target domain. On the other hand, anchors enable agents to learn to compromise as shown in Fig. 2(c). The evolution of the pendulum angle is shown in Fig. 3. Additional data for more agents and for agents trained with different algorithms are available in Section VIII.
Figure 3: The sin of the inverted pendulum’s angle $\alpha$ over time starting from $\alpha=90^{\circ}$ for a DDPG agent trained initially on Domain A as per Fig. 2(a), and then adapted to Domain B with Domain A anchors as per Fig. 2(c). Note that initial training allows the agent to stably balance the pendulum to approximately $10^{\circ}$ while adaptation results in the agents driving the pendulum to approximately $0^{\circ}$.
IV-B Catastrophic Forgetting in Reacher
In Fig. 4, we showcase the severity of catastrophic forgetting when fine-tuning agents naively. This issue is especially grave on the reacher task since the target domain has exactly the same dynamics as the source domain, only differing in setpoint locations such that the ones in the target domain are a strict subset of the ones in the source. The unstable behavior of the naively fine-tuned agent is a result of the agent overfitting to the target domain and forgetting how to solve the task in the source domain.
Figure 4: Evolution of Reacher agents on the source domain over time (arrows showing directionality). We contrast a naively fine-tuned agent with one that has been anchored. The plots show how unstable the naively tuned agent is, while the anchored agent is able to maintain stability.
IV-C Preventing Catastrophic Forgetting on Gymnasium
We perform experiments on three benchmark Gymnasium environments, namely Reacher-v4, LunarLanderContinuous-v2, and Pendulum-v0. For each environment, we vary some parameters as described in Table I to test the impact of fine-tuning when a domain shift occurs.
Environment Source Domain Target Domain
Pendulum-v0 gravity 13m/s2 gravity 4m/s2
Reacher-v4 goal within 0.2m goal within 0.1m
LunarLanderContinuous-v2 gravity 10m/s2 gravity 2m/s2
TABLE I: Source and target domains for each environment.
[…]
We then show that anchors can mitigate this issue accross multiple environments with reality-gaps both in dynamics and observation distributions in Fig. 5.
[…]
V Anchor Critics in Reality
To address the challenges of consistent and safe live-adaptation with reinforcement learning in reality, we evaluate the effectiveness of anchor critics through a series of experiments:
  1. 1.
    Leveraging SwaNNL, we evaluate how live-adapting with and without anchors affects safety, smoothness, and performance on the real drone system in Sections VI-B and VI-C. The novel infrastructure offered by the SwaNNL enabled us to develop and test practical ways to adapt agents while a quadrotor is in flight, demonstrating how anchors make live adaptation safer on real drones.
  2. 2.
    We offer theoretical arguments motivating our method in Section III.
  3. 3.
    Finally, we discuss known limitations with respect to drone control and the general cost of using anchors in Section IX.
[…]
V-A Engineering SwaNN Lake for Live NN updates
Figure 6: SwaNNL overview for live adaptation. 1) Observations received as a result of policy invocation are packaged and 2) sent to the ground-station where they are unpacked and fed into an RL algorithm. 3) The updated policy is packaged with integrity check markers and 4) transmitted to the robot. 5) Upon verification of the received policy graph, the new graph is copied to system memory and 6) interpreted by TFLite for invocation.
In this work, we investigate strategies for RL-based live control adaptation on real robots. As such, two important sub-goals are formulated as follows. First, we needed to allow BetaFlight [34]-compatible firmware to update its NN controller(s) without interrupting the control loop. As BetaFlight is a standard across various robotic platforms, this allows our framework to support a large range of robots, such as blimps, boats, helicopters, racing drones, and more. Second, we needed to be able adapt the agents live without crashing. This necessitated preventing catastrophic forgetting.
We thus decided to evolve Neuroflight [10, 11], initially developed by Koch et al. for allowing the use of NNs for high-performance quadrotor control. Neuroflight extends the open-source Betaflight [34] flight-control firmware stack allowing NN models to be compiled and embedded into the control firmware. We evolved the Neuroflight framework such that modifications to the employed NN controller(s) can be carried out without interrupting the control loop. To do so, we re-designed the firmware so that we could swap out binarized NN graphs with minimal overhead rather than having them directly compiled into the firmware. In particular, it is possible to both (i) update the NN controller weights and (ii) swap-in graphs with different NN architectures. Our novel firmware stack outlined in Fig. 6, is open-sourced [30]. Our architecture allows inference to be performed onboard the micro-controller, but training onboard is prohibitively expensive, given compute and storage limitations of low-power embedded platforms. We thus introduce a point-to-point communication channel with an external ground station, to mitigate this issue. Most importantly, the ground station can be used as a training platform to adapt controllers based on data captured in flight. SwaNNFlight further integrates with wireless transceiver modules to (i) wirelessly transmit flight data and (ii) accept new NN models to switch to mid-flight.
VI TRAINING FLIGHT CONTROLLERS WITH SwaNNL
Prior work employing the Neuroflight training framework trained RL-based controllers using TF1 [10, 11, 35, 36]. Attempting to deploy these RL-based controllers using our updated stack with TFLite’s interpreter resulted in the control firmware crashing during invocation, due to the operations used by the translated TFv1 NNs. Agents built with TF2, conversely, could be successfully deployed and invoked. This compelled us to update our training stack to TF2.
Transitioning to TF2 introduced a problem for the training pipeline. Prior works [36, 35, 11] train flight-control agents with the OpenAI Baselines [37] implementation of PPO [19]. While reproducible with the software packages and versions reported by the authors, the performance observed with PPO agents in TF1 was not achievable by any of several tested TF2 implementations of PPO, including TF-agents [38], stable-baselines [39] etc., despite using identical network architectures and performing a full hyperparameter search. Algorithms like PPO and TRPO [18] are known to be sensitive to (and even brittle under) implementation specifics [40, 41, 42], prompting a search for alternatives.
Viable flight-control policies were trainable using a TF2 translation of OpenAI’s Spinning Up [43] PyTorch implementation of DDPG [31]. While vanilla DDPG presented with similar oscillatory behavior identified in the work on CAPS  [36], introducing CAPS regularization to DDPG enabled the training of viable agents in simulation. We found, however, that the performance of DDPG on the control problem at hand exhibited relatively high variance: not all the trained agents were safely flyable. This prompted us to explore alternative formulations which improve upon CAPS defined in the next section.
VI-A SwaNNFlight System Transceiving Metrics
Wireless communication is handled by Digi XBee® ZigBee®-PRO radio-frequency modules. The drone communicates observation data to an XBee through a UART port, while the ground station uses an XBee to send updated network graphs back to the drone. This addition represents the only hardware change, with negligible impact on weight and power consumption.
To ensure data integrity, we implemented a three-phase handshake protocol with cyclic redundancy checks (CRC). Upon verification, the drone atomically swaps to the new graph at the next control cycle. Buffered data is automatically chunked into CRC-validated packets, supporting multiple transmission rates.
Using the MATEK-F722 controller with ZigBee®-PRO modules, we measured transmission and switching times for neural networks (NNs). Our control networks use two hidden layers with 32 neurons each. At a baudrate of 115200, sending an 8mb NN takes $\approx{11}$ seconds, with receiving handled in parallel to flight control. Switching to a new NN controller takes $\approx{134}$ ms.
During flight, the drone transmits 59-byte state observations at a rate of 244 per second to the ground station.
VI-B Catastrophic Forgetting and Safety Concerns in Adaptation without Anchors
Before testing adaptation in live, unconstrained flight, we first conducted a series of tests in a controlled lab environment where the drone was attached to a light tether to restrain it in the event of a catastrophic loss of control, without being a hindrance to flight. These tests consisted of fine-tuning a well-trained controller from simulation (the source domain) on the real platform (the target domain) while giving it typically small control targets ($<50$ deg/s) and occasionally requesting control in excess of that ($>100$ deg/s). Without anchors preserving agents’ performance on larger control inputs—like those experienced during the simulated training—the live agents were observed to forget how to maintain tracking when receiving large inputs, though sometimes regaining control, and often losing it again, as seen in Fig. 7. This also results in a sudden loss of control where errors would quickly snowball and motor actuation rises sharply, risking damage to the drone. Fig. 8 demonstrates how, when commanding attitudes not seen for a while during live-flight adaptation, agents could very quickly lose control (observed from MAEs increasing quickly and exponentially) – which we attributed to overfitting over the relative homogeneity of the live-flight training data akin to the reacher environment we tested. Anchors help avoid this problem.
Figure 7: Shown here is how MAEs and smoothness evolve with adaptation steps, with and without anchors. In both cases smoothness improves, however, without anchors, tracking errors spike often as catastrophic forgetting occurs. With anchors, error is bounded and remains in controllable range.
Figure 8: A snapshot of instability of the drone during an adaptation run without anchors. Observe how error increases exponentially as the drone controller becomes unstable.
(a) Linear Composition
(b) Multiplicative Composition
Figure 9: When comparing multiplicative composition of optimization criteria, as described in Section VII, against linear composition, we observe that the multiplicative composition quickly and significantly reduces learning variance while achieving comparable or better performance on the tracking error, motor acceleration, and motor actuation objectives.
VI-C Live Adaptation During Unconstrained Flight
TABLE II: Errors and power consumption in flight before and after 15 steps of live adaptation
Before Adaptation After Adaptation
Amperage $\downarrow$ $13.7\pm 8.47$ $7.24\pm 3.97$
Mean Average Error $\downarrow$ $12.55\pm 12.22$ $14.13\pm 5.21$
Smoothness ×$\times 10^{4}$downarrow $12.6\pm 0.98$ $5.85\pm 0.96$
Figure 10: Comparisons of motor usage during flight and corresponding FFTs demonstrate the improvement in smoothness after adaptation, as noted from the reduction in high-frequency components on the FFT amplitude plot. While the motor usage plot samples a single trajectory for clarity, the FFTs are averaged over multiple adaptations to demonstrate repeatability.
We use two key performance metrics in policy evaluation. (1) Tracking accuracy is measured using the Mean Absolute Error (MAE). (2) Smoothness is computed by taking the Fast Fourier Transform (FFT) of the actuation commands, and then computing $Sm$ as defined by Equation 4 in [36] (lower is better). Shown in Fig. 10 is a comparison of control signals before and after 15 adaptation steps. Motor usage is reduced significantly, resulting in improved control smoothness and almost a 50% reduction in power consumption, with comparable MAEs to agents trained in simulation. This is further demonstrated in Table II, which examines how MAE, smoothness, and current draw change through adaptation.
VII Multiplicative Composition of Optimization Objectives
Multiplicative composition of multi-objective rewards through a geometric mean, with each reward in [0,1], has been shown to improve training performance and reduce policy variance compared to linear composition [35]. This approach, analogous to a smooth AND operator, incentivizes policies to satisfy all defined objectives. We apply this concept to our optimization objectives and regularization terms, improving upon the linear composition used in CAPS [36].
For an RL policy $\pi_{\theta}$, the DDPG policy optimization criteria under our multiplicative version of the CAPS regularization scheme is:
$J_{\pi_{\theta}}=\sqrt[4]{Q_{\pi_{\theta}}\left(\frac{w_{T}}{w_{T}+L_{T}}\right)\left(\frac{w_{S}}{w_{S}+L_{S}}\right)\left(\frac{w_{A}}{w_{A}+\pi_{\theta}^{:-1}}\right)}.$ (12)
where: $L_{T}$ and $L_{S}$ are temporal and spatial smoothness regularization terms, $Q_{\pi}$ is the critic estimating Q-values of state-action pairs, $\pi_{\theta}^{:-1}$ represents the policy network output prior to the final $\tanh$ activation, $w_{T}$, $w_{S}$, and $w_{A}$ are penalty thresholds for $L_{T}$, $L_{S}$ and $\pi_{\theta}^{:-1}$ respectively.
[…]
VII-A Ablation Study on Multiplicative Composition
To motivate the use of multiplicative composition in our optimization criteria, we trained controllers in simulation using CAPS-like [36] linear loss composition against multiplicative loss composition, shown in Equation 12. We conducted extensive hyperparameter search with different weighting parameters, both for the CAPS regularization terms as well as the training rewards. Even with our best set of hyperparameters, however, variance in trained performance remained high. We show in Fig. 9 that we achieved higher performance with less tuning and significantly reduced variance using multiplicative composition. The results presented in the figure compare 6 independently trained agents for both linear and multiplicative composition. Weights shown in Equation 12 are chosen so that undesirable performance on any criterion would have values near 0.
TABLE III: Performance comparison of RL algorithms. Training: 300,000 environment invocations, 5 tests per hyperparameter sample. Rewards computed for 10+ samples per algorithm. Step times: Ubuntu v18.04, Intel Core i7-7700 CPU (no GPU optimization).
Algorithm DDPG×$\times$ DDPG PPO (TF2) SAC TD3
Reward $\scriptstyle 0.89\pm 0.033$ $\scriptstyle 0.38\pm 0.13$ $\scriptstyle 0.36\pm 0.11$ $\scriptstyle 0.40\pm 0.12$ $\scriptstyle 0.37\pm 0.11$
Max Rw 0.930.93$\scriptstyle$ 0.930.93 0.590.59$\scriptstyle$ 0.590.59 0.610.61$\scriptstyle$ 0.610.61 0.630.63$\scriptstyle$ 0.630.63 0.570.57$\scriptstyle$ 0.570.57
Steps (s) $\scriptstyle 9.6\times 10^{-3}$ $\scriptstyle 9.6\times 10^{-3}$ $\scriptstyle 3.6\times 10^{-3}$ $\scriptstyle 1.44\times 10^{-2}$ $\scriptstyle 1.2\times 10^{-2}$
VIII Hyper-parameter search
We conducted a hyper-parameter search and a correlation factor analysis using 4 different prominent RL algorithms – PPO, DDPG, SAC and TD3 – for the quadrotor attitude control problem mainly considered in this paper. The list of tested hyper-parameters and their corresponding correlations to the training reward are presented in Tables IVVVI and VII respectively. Our hyper-parameter choices were informed by common values used to achieve the best performance across various benchmark Gymnasium environments [44].
Table III meanwhile compares the training performances of each of the algorithms, as well as DDPG×$\times$ which we introduce in this work, after 300000 training steps (corresponding to a the time by which DDPG×$\times$ typically starts to plateau), but on a slightly modified version of the problem – where agents are only required to minimize tracking error, without worrying about other characteristics such as control smoothness. This, we found, was typically a good starting point for tuning algorithms’ hyper-parameters and to determine if they are capable of addressing the underlying control problem. As the data demonstrates, all the tested algorithms perform significantly worse than DDPG×$\times$.
While the relative performance metrics appear similar across DDPG, PPO, SAC and TD3, there are critical differences in the quality of the behaviors being learned, which in turn impacts the ability of the algorithms to further improve agents’ performances. As shown in Fig. 11, the tested TF2 implementation of PPO does not behave similarly to the modified OpenAI Baselines [37] TF1 version used in previous work. [36, 35, 11]. Where the PPO based on Baselines learns to track the control signals well, albeit with high variance in the motor actuation (since smoothness is not being trained for), the TF2 implementation saturates the motor commands at the maximum, thus effectively doing nothing to control the drone’s attitude. This still generates some reward, since the training environment predominantly consists of smaller (but more frequently varying) control angular velocities. The problem with this, however, is that this behavior seems to be a local optima that is difficult for the algorithms to recover from to further improve performance.
When considering alternatives to PPO, we experimented with DDPG, SAC and TD3. As shown in Fig. 12, we observed that TD3 would often fall into similar trappings as PPO, with the motor outputs becoming saturated at maximum actuation. While DDPG was slower to learn (often performing worse than SAC), crucially, it did not saturate and would in fact consistently try to maintain control through more nuanced use of the motors – an extreme but illustrative case is shown in Fig. 12. Given our results, while SAC would have been preferable as a starting point based on the typical rewards achieved, the fact remained that training SAC was approximately 1.5 times slower than DDPG for relatively marginal gains, making it a less attractive option for iterating. We therefore chose DDPG to experiment on our proposed multiplicative policy optimization. In our tests, agents which achieve an average reward of more 0.8 were typically responsive enough to control in real flight.
TABLE IV: TF2 PPO paramater/reward correlations
Parameter Correlation Tested Values
Clip ratio 0.098 [0.0, 0.05, 0.1, 0.2]
$\gamma$0.100 [0.8, 0.9, 0.95, 0.99]
Policy LR 0.679 [1e-4, 5e-4, 1e-3, 3e-3, 5e-3]
Value LR 0.255 [1e-4, 5e-4, 1e-3, 3e-3, 5e-3]
lam 0.056 [0.9, 0.97, 0.99, 0.995]
TABLE V: DDPG paramater/reward correlations
Parameter Correlation Tested Values
Buffer size -0.156 [1e5, 5e5, 1e6, 5e6]
$\gamma$0.161 [0.8, 0.9, 0.95, 0.99]
Polyak 0.203 [0.5, 0.9, 0.95, 0.99, 0.995]
Policy LR -0.211 [1e-4, 5e-4, 1e-3, 3e-3, 5e-3]
Critic LR -0.676 [1e-4, 5e-4, 1e-3, 3e-3, 5e-3]
Batch size -0.142 [50, 100, 200, 400]
Action noise -0.691 [0.01, 0.05, 0.1, 0.2]
Figure 11: A snapshot of PPO and DDPG×$\times$ agents after 300,000 training steps acting in their training environment. Attitude control elements are split into Yaw, Pitch, and Roll in the left column. On the right are the outputs generated by the agents while attempting to follow the control inputs. Without smoothness considerations, within 300,000 steps, Baselines PPO performs very well on tracking while outputting actions with high variability (as is shown in previous works [36], [35]). Conversely, the TF2 implementation of PPO fails to learn useful behavior across hyper-parameters. DDPG×$\times$, which we introduced in this work, is able to reasonably follow the target while maintaining smooth outputs and continues to improve as training progresses.
TABLE VI: SAC paramater/reward correlations
Parameter Correlation Tested Values
Buffer size -0.088 [1e5, 5e5, 1e6, 5e6]
$\gamma$ 0.272 [0.8, 0.9, 0.95, 0.99]
Polyak 0.102 [0.5, 0.9, 0.95, 0.99, 0.995]
LR -0.256 [1e-4, 5e-4, 1e-3, 3e-3, 5e-3]
$\alpha$ 0.356 [0.01, 0.05, 0.1, 0.2]
Batch size 0.866 [50, 100, 200, 400]
TABLE VII: TD3 paramater/reward correlations
Parameter Correlation Tested Values
Buffer size 0.516 [1e5, 5e5, 1e6, 5e6]
$\gamma$ 0.479 [0.8, 0.9, 0.95, 0.99]
Polyak 0.007 [0.5, 0.9, 0.95, 0.99, 0.995]
Policy LR -0.430 [1e-4, 5e-4, 1e-3, 3e-3, 5e-3]
Critic LR 0.023 [1e-4, 5e-4, 1e-3, 3e-3, 5e-3]
Batch size 0.118 [50, 100, 200, 400]
Act noise 0.313 [0.01, 0.05, 0.1, 0.2]
Figure 12: For the three DDPG-based algorithms, we observe clear dissimilarities in the quality of the policies learned, based on input tracking and motor usage. Entropy regularization employed in SAC allows trained agents to achieve relatively reasonable performance on both tracking and motor usage, but the algorithm was noted to not improve further with more training, likely due to the same regularization that initially helps preventing large responses, thus preventing the algorithm from learning to track more aggressive maneuvers. TD3 behaves similarly to the TF2 PPO, in that it tends to saturate at various local optima, outputting extreme values. DDPG, on the other hand, demonstrates some tracking capabilities can varies its control output drastically when close to the setpoint, similar in behavior to Baselines PPO. Given these three candidate algorithms and their relative benefits and drawbacks, we chose to modify DDPG as it runs faster than SAC and was capable of achieving good tracking performance. While DDPG itself was not consistently reliable, DDPG×$\times$, with our proposed improvements, was repeatable and controllable.
IX Known Limitations
While anchors clearly reduce power consumption without resulting in catastrophic forgetting, the adapted agents were not observed to improve their tracking performance significantly, even though this was one of the key learning objectives of the policy optimization. It is as yet unclear as to whether this is simply a limitation of the DDPG algorithm (or our specific implementation), or a matter of how the regularization was tuned. In cases where domain gaps are large, the very stability offered by the anchor critics may significantly compromise fine-tuning of behavior and could result in policies that are ill-suited for both the initial and transferred domain. Finally, our usage of a saved replay buffer from simulation to evaluate anchor critics limits the ability to adapt anchors to new policies, as a divergent enough policy will make anchors an unsuitable q-value for the source domain. We suspect that keeping the simulation in the loop might allow for anchors to allow for continued learning long-term.